Find how much money needs to be deposited now into an account to obtain $7700 in 15 years if the interest rate is 8% per year compounded monthly
Answer by Boreal(15235)

This is P=Po(1+(r/t))^15
7700=Po(1+(.08/12))^180, which is number of months in 15 years
7700=Po(3.307) but don't round on yet using a calculator.
Po=$2328.45